OpenStudy (firejay5):

Solve each equation: Please be clear and make on sense on solving equation with/without instructions: 5(6 - 4X) = X + 21

OpenStudy (lgbasallote):

first, distribute 5 into (6-4x) you're familiar with the distributive property yes?

OpenStudy (firejay5):

yes

OpenStudy (lgbasallote):

so what is the result when you distribute 5 into (6-4x)

OpenStudy (firejay5):

30 - 20X

OpenStudy (lgbasallote):

good... so the whole equation will become 30 - 20x = x+ 21

OpenStudy (lgbasallote):

now add 20x to both sides so that all the x are just on one side

OpenStudy (firejay5):

yes so it's 30 =21X + 21

OpenStudy (lgbasallote):

right. now subtract 21 from both sides

OpenStudy (firejay5):

9 = 21X

OpenStudy (lgbasallote):

and lastly, divide both sides by 21

OpenStudy (firejay5):

3/7

OpenStudy (lgbasallote):

right!

OpenStudy (firejay5):

is it X = 3/7

OpenStudy (firejay5):

is that the answer

OpenStudy (lgbasallote):

yes x = 3/7 is the final answer

OpenStudy (firejay5):

can you please show how it would it look like put in for X with the answers

OpenStudy (lgbasallote):

what do you mean?

OpenStudy (firejay5):

enter 3/7 in for X

OpenStudy (lgbasallote):

do you mean for checking? 5(6 - 4(3/7) ) = 3/7 + 21 \[\implies 5(6 - \frac{12}{7} ) = \frac 37 + \frac{147}{7}\] \[\implies 5(\frac{42}{7} - \frac{12}7) = \frac{150}{7}\] \[\implies 5(\frac{30}{7}) = \frac{150}{7}\] \[\implies \frac{150}{7} = \frac{150}7 \checkmark\]

OpenStudy (lgbasallote):

like that?

OpenStudy (firejay5):

ya thanks! I could figure it out how it worked

OpenStudy (firejay5):

How did you get 147/7

OpenStudy (lgbasallote):

to make 21/1 into a fraction with denominator 7, first divide 7 by 1. the answer is still 7. then, multiply 7 to 21 which is 147

OpenStudy (firejay5):

okay gotcha

OpenStudy (firejay5):

how did you get 42/7

OpenStudy (lgbasallote):

6/1 divide 7 by 1 hat becomes 7 then multiply to 6..42

OpenStudy (firejay5):

hello dude what's going on #lgbasallote